OpenStudy (arbyj):

Someone help me solve; 10x-15=10x+4

OpenStudy (anonymous):

so first you get all of the same term on one side so you you get (10x-10x) = (4+15) which is 0=19 which is not true so there are no solutions. (your answer is "no solutions exists").

OpenStudy (anonymous):

10x-15=10x+4 10x-10x-15=4 0=19, but this is not true, so no solution